I have tried to find the equation in standard form 10 times and I get x-2y=15 however the answer is x-2y=-15 and it's driving me crazy. The original point and slope is (-3,6);m=1/2 I used the formula y-y1=m(x-x1). please help
Answer by Alan3354(69443)

The original point and slope is (-3,6);m=1/2 I used the formula y-y1=m(x-x1)
--------------
Slope m = 1/2 thru the point (-3,6)
Use y = mx + b and the point
6 = (1/2)*(-3) + b
b = 15/2
--> y = x/2 + 15/2
2y = x + 15
x - 2y = -15
---------------
Using y-y1=m(x-x1)
y - 6 = (1/2)*(x+3)
2y - 12 = x + 3
x - 2y = -15